X, a non-Government employee, receives Rs. 3,06,000 as leave salary at the time of retirement on February 20, 2021. On the basis of the following information, determine the amount of taxable leave salary: Basic pay: Rs. 18,000 per month since 2013; duration of service: 32 years; leave at the credit of X at the time of retirement: 17 months ; entitlement of leave salary: 45 days' salary for every year of service and leave availed while in service: 31 months.


1
Expert's answer
2021-10-05T09:20:43-0400

Solution:

First, derive how much tax will be exempted and then deduct from the amount paid to get the taxable leave salary.

The tax exemption would be a minimum of the below 4 points:

·                    Amount received as leave encashment – Rs 306,000

·                    Maximum cap as stated by the government – Rs 300,000

·                    Last 10 months average basic salary & dearness allowance before leaving the job – Rs 180,000 (Rs 18,000 "\times" 10)

·                    Cash equivalent of the leave balance, subject to a maximum of 30 days for each completed year of service – Rs 306,000 (as per calculation below)


Leaves eligible for encashment (as per above rule) = 17 months

Cash equivalent = 17 "\times" 18,000 = Rs 306,000

Tax exemption = Rs 180,000

Taxable leave salary = Rs 306,000 – Rs 180,000 = Rs 126,000